1) Property of inverse functions If f(x): (-1,4 ] → [3,7) → this means f(x) is a function with domain (-1,4 ) and range (3,7] answer all please!
Answer #1

2 Range of f que [Let ► fl): 1-1,4] → [3,7] Domain of f = Range of par = [1,4] Domain of p- [3,7] f(x) : [3,7) → [-1,4] R>a)3 dy f(x)=n >> y = 23 [Let y=f(x)] y >> f-ly)2 ys since, x = f(y)] >>4-() 27 23 20 37a) f(x) = sinx by se z cesy AY an →

If you have any problem while understanding any part, Leave a comment. I will try my best to help.